OpenStudy (anonymous):

A multiple choice test has 8 questions. Each question has 4 possible answers. Find the probability that a students who guesses the answer for each question will pass the test. Show all steps

OpenStudy (thesecret20111):

Ok use the probability formula P=number of desired outcomes/number of total outcomes

OpenStudy (anonymous):

so would it be \[\frac{ 4 }{ 32 }\] ?

OpenStudy (thesecret20111):

Wait it's been ages since I've taken and maths lessons in probability.. You might need to use a different formula to make it easier.

OpenStudy (thesecret20111):

Are you familiar with Binomial Probability Formula?

OpenStudy (thesecret20111):

It goes like this: P(k success in n trials) =\[\left(\begin{matrix}n \\ k\end{matrix}\right)p ^{k}q ^{n-k}\]

OpenStudy (thesecret20111):

n = number of trials k = number of successes n – k = number of failures p = probability of success in one trial q = 1 – p = probability of failure in one trial

OpenStudy (thesecret20111):

For this question then n = 8 k = 8 n – k = 0 p = 0.25 = probability of guessing the correct answer on a question q = 0.75 = probability of guessing the wrong answer on a question

OpenStudy (thesecret20111):

Now plug them in and use a calculator.